OpenStudy (anonymous):

find (a) the GCF and (b) the LCM of the following monomial 9x^2y and 36xy^3

OpenStudy (anonymous):

help plz can u tell me how to find them

OpenStudy (cwrw238):

the GCF - the highest common factor of 2 or more numbers is the highest number which divides exactly into them what do u think the GCF of 9 and 36 is?

OpenStudy (cwrw238):

similarly for variables what is GCF of x and x^2? - it is x

OpenStudy (anonymous):

9 would be the highest common factor right

OpenStudy (cwrw238):

taks 2 other numbers - 2 and 6 - the GCF is 2 because 2 divides into 2 and 6 and its the greatest number able to do this

OpenStudy (cwrw238):

yes - 9 is correct

OpenStudy (cwrw238):

highest - greatest same thing in US they say GCF whereas in UK its HCF

OpenStudy (anonymous):

but why is it - ? it would be 9 x no ?

OpenStudy (cwrw238):

9 is GCF of number bit only x is GCF of x and x^2 and you now need to find GCF of y and y^3 which is what?

OpenStudy (anonymous):

its y

OpenStudy (anonymous):

so the final answer would be 9xy because these are the gcf of all three

OpenStudy (cwrw238):

dead right so the GCF of the two expressions is 9xy

OpenStudy (anonymous):

wow thanks alot! u explained it very well i realli appreciate it :)

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

:)

OpenStudy (cwrw238):

LCM = lowest common multiple eg LCM of 4 and 8 is 8 lcm of 6 and 9 is 18

OpenStudy (cwrw238):

for 9 and 36 it is 36

OpenStudy (cwrw238):

a sure way of finding lcm - especially for large numbers is to find the prime factors first eg 6 and 9: 6 = 2 * 3 9 = 3 * 3 as there is a match of one pair of 3's in the two factors you only use 3 once when working out lcm- you also include all other factors so we get 2 *3 * 3 = 18

OpenStudy (cwrw238):

for x and x^2 lcm = x^2 for y and y^3 lcm = what

OpenStudy (anonymous):

so 3 *3 = 9 6*6= 36 im confused

OpenStudy (cwrw238):

there is no need to use the factor method for 9 and 36 because 9 divides into 36 exactly so the lowest common multiple must be 36 but you haven't got the PRIME factors for 36 : - 6 not prime number 9 = 3 * 3 36 = 2 * 2 * 3 * 3 there's 2 pairs of matching 3's so we need 3 * 3 so lcm = 2*2 *3*3 = 36 i hope i haven't complicated things too much if you have numbers like 2 and 8, or 3 and 9, 12 and 48 where one number divides exactly into the other the lcm is of course the largest of the 2 numbers

OpenStudy (anonymous):

i did an LCM of 28 35 and 42 and i got 7

OpenStudy (anonymous):

thank u very much for taking the time to make me understand